OpenStudy (anonymous):

Find a cubic function with the given zeros. Square root of six., - Square root of six., -3 f(x) = x3 - 3x2 - 6x - 18 f(x) = x3 + 3x2 - 6x - 18 f(x) = x3 + 3x2 + 6x - 18 f(x) = x3 + 3x2 - 6x + 18

OpenStudy (triciaal):

(x +6)(x-6)(x+3) = 0

OpenStudy (anonymous):

i have x^3+3x^2-36-108=0

OpenStudy (triciaal):

you have to multiply each term someone else was also responding. i guess he changed his mind i will do it now

OpenStudy (triciaal):

you forgot your x then factor the 6

OpenStudy (triciaal):

x^3 + 3x^2-36x-108

OpenStudy (anonymous):

\[(x-\sqrt{6})(x+\sqrt{6})(x+3)=0\] I think you can expand from here.

OpenStudy (anonymous):

how did you get it?

OpenStudy (triciaal):

sorry i made a mistake it should be as above rt 6

OpenStudy (triciaal):

(x-rt6)(x +rt6)(x+3) recognise the difference of 2 squares? (x^2-6)(x+3) expaND X^3-6X+3X^2-18 Rearrange to put in order of degree x^3 +3x^2 -6x-18

OpenStudy (triciaal):

sorry recognize